rssLink RSS for all categories
 
icon_blue
icon_green
icon_green
icon_orange
icon_red
icon_green
icon_green
icon_orange
icon_red
icon_red
icon_green
icon_green
icon_green
icon_blue
icon_orange
icon_red
icon_green
icon_blue
icon_red
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_orange
icon_green
icon_green
icon_red
icon_green
icon_blue
icon_green
icon_green
 

FS#33932 — Fibre BHS <> Vaudreuil

Attached to Project— Reseau Internet et Baies
Incident
Beauharnois
CLOSED
100%
Nous avons une coupure fibre confirmée entre le DC de BHS et le site de Vaudreuil. Ceci impact nos circuits optiques vers Newark qui sont passés automatiquement en protection mais aussi les circuits BHS <> Ashburn, BHS <> Toronto et BHS <> Chicago.

We have got a fiber cut confirmed between the Beauharnois datacenter and Vaudreuil. This impact our optical circuits towards Ashburn, Toronto and Chicago
Date:  Thursday, 06 September 2018, 19:26PM
Reason for closing:  Done
Comment by OVH - Wednesday, 05 September 2018, 07:20AM

According to the latest update received from our on-site teams. They just finished the trace tests on the first point. They go to the second point to complete the tests to accurately determine the breakpoint. We will keep you informed of developments as soon as we receive new information.


Comment by OVH - Wednesday, 05 September 2018, 17:00PM

The fibercut is localised on a bridge at Valleyfield, QC.
There are about 400 fibers down.
The provider needs to rolled out 3km of new cable to repair.
They just had authorizations from MTQ to begin the work on the bridge.

ETR is 9:00 PM Local Time.


Comment by OVH - Thursday, 06 September 2018, 03:59AM

Repair works are still ongoing. Initial impact seems to have been caused by vandalism. A new cable is being installed on 2 sections. Cable installation has been completed on first section. Second section is currently in progress.


A new ETR has been provided : fiber pairs continuity will be restored from 12:30 AM local time


Comment by OVH - Thursday, 06 September 2018, 07:42AM

The installation of the second section continues with several difficulty.

Modification of the ETR, a gradual restoration of services is now expected from 5:00am local time


Comment by OVH - Thursday, 06 September 2018, 14:13PM

All circuits are now up.

We are waiting the confirmation that all works are finished